# A Noisy Approach to Intrinsically Mixed-State Topological Order

### v2

@@ III.4 Three-Fermion from $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ Toric code
As a final remark, finite temperature mixed-states also provide simple instances of our general framework. For example, consider the $D=2,3,4$ Toric code at finite temperature [51, 12]. For $D=2$, any finite-T state corresponds to both the $e,m$ anyons being condensed: the resulting state hosts no deconfined anyons and is hence trivial. Now, in $D=3$, the $e$ charges proliferate at any non-zero temperature, but below a critical $T_{c}$, the flux-loops of the 3D Toric code remain deconfined but are now transparent. This corresponds to a TSSC that does not encode any logical qubits in its logical subsystem but still has a non-trivial classical memory due to the transparent loops. Finally, for the 4D Toric code which has only loop-like excitations, there exist two critical temperatures: below the first, none of the excitations proliferate and the finite-T mixed-state is a TSSC that is equivalent to a topological stabilizer code (i.e., its gauge group is proportional to the stabilizer group). Hence, it represents finite-T topological order. Above the first, but below the second critical temperature, only one of the loop excitations proliferate and one obtains a TSSC with a classical memory. Above the second critical temperature, all anyons are condensed and the mixed-state is topologically trivial. Thus, prior results on topological order at finite temperature are straightforwardly incorporated into our general framework. We note that an infinite temperature state with quantum memory based on a subsystem code was previously proposed in Ref. [52], and in our framework, constitutes an imTO.
In general, the map from decohered density matrices to TSSCs conveniently allows one to use results regarding the latter to obtain a partial classification of the former. In particular, since Ref. [36] showed that any (non-modular) Abelian anyon theory can be realized by a TSSC, it immediately provides a partial classification of imTOs in terms of Abelian anyon theories.
@@ IV Decoherence as gauging out in general UMTCs
Thus, we can now define gauging out anyons in the same way as before, but in a more general context: starting with a UMTC $\mathcal{C}$ and a proper subset of anyons $\hat{\mathcal{A}}$ to be gauged out, the resulting anyon theory (the code space of the decohered theory) is given by those anyons in $\mathcal{C}$ which braid trivially with those in $\hat{\mathcal{A}}$. Moreover, if any anyons in $\hat{\mathcal{A}}$ are transparent in $\hat{\mathcal{A}}$, they remain transparent in $\mathcal{A}$, which will generically be a braided fusion category (without the modularity restriction). Formally, given a UMTC $\mathcal{C}$ and a proper subset of objects (anyons) $\hat{\mathcal{A}}$ (i.e., a full subcategory of $\mathcal{C}$), the anyon theory $\mathcal{A}$ that results upon gauging out $\hat{\mathcal{A}}$ is given by the centralizer $C_{C}(\hat{\mathcal{A}})$ of $\hat{\mathcal{A}}$ in $\mathcal{C}$: $\mathcal{A}\equiv\{x\in\mathcal{C}|\mathcal{B}_{\theta}(x,y)=1\,\forall y\in\hat{\mathcal{A}}\}$, which is a braided fusion category (see Ref. [57]). One could in principle then generate another braided fusion category by gauging out anyons from $\mathcal{A}$ and generate a cascade of imTOs by iteratively gauging out anyons.
We believe that this picture for obtaining braided fusion categories from parent UMTCs falls squarely within the general class of mixed TQFTs proposed by Zini and Wang in Ref. [58], but where the input to the parent 2+1D Turaev-Viro (TV) type TQFT is always modular. In our context, this restriction is physically motivated since we take as input the ground state of a local Hamiltonian (so the anyon theory is a UMTC) and then subject it to local noise. In fact, note that the doubled semion example we previously considered is one where the imTO we obtained belongs to the class of mixed TQFTs considered in Ref. [58]. Specifically, in that case the input was the doubled semion UMTC and the output imTO was the chiral semion UMTC, where we can view the latter as the Reshetikhin-Turaev (RT) TQFT of the former: this supports our claim that braided fusion categories provide a partial classification of imTOs. More generally, if the parent theory is some doubled Chern-Simons theory (which have gapped boundaries to vacuum), one might expect that local error channels will lead to the underlying chiral Chern-Simons TO in the decohered mixed-state—we investigate this in a forthcoming work. For now, we consider some simple examples that go beyond Abelian anyon theories to show the generality of our framework.
@@ IV.1 $\mathbb{Z}_{2}^{(1)}$ TSSC from Chiral Ising UMTC
A physical Hamiltonian that supports a phase with chiral Ising TO is furnished by the Kitaev honeycomb model [46]. Recently, Ref. [59, 27] considered the honeycomb model in the presence of local $\psi$ errors, which are generated by a local error channel. We will not dwell on details of the lattice model but directly reproduce the results using our general framework. Specifically, we wish to gauge out the $\psi$ fermion: since $\sigma$ braids non-trivially with $\psi$, the resulting anyon theory describing the decohered state is simply given by $\{1,\psi\}$, which does not encode any quantum memory but still yields a classical memory and retains a well-defined fermionic excitation. This is in precise agreement with the numerical results of Refs. [59, 27], which also verified the stability of this intrinsically mixed-state TO, which we expect on general grounds from the unbroken strong anomalous 1-form symmetry of $\rho_{\mathcal{E}}$.
@@ IV.3 Non-modular imTOs from Doubled $SU(2)_{k}$ UMTC
As a final example, we can consider doubled SU(2)${}_{k}$ string-net models, whose lattice models and short-string operators are given in Ref. [54]. Anyons in this theory are labeled by pairs $(j_{1},j_{2})$ where $j=0,1/2,1,\dots,k/2$. Let us now subject a ground state of this model to local errors that incoherently proliferate the anyon $(k/2,k/2)$ (which is a boson). In order to read off the resulting imTO in the decohered density matrix, after $(k/2,k/2)$ has been gauged out, we require the braiding relations of this theory. In particular, the braiding between an anyon $(j_{1},j_{2})$ and $(k/2,k/2)$ is given by $\mathcal{B}_{\theta}((j_{1},j_{2}),(k_{1},k_{2}))=(-1)^{2(j_{1}+j_{2})}$. Thus, the resulting imTO is characterized by the anyon theory $\mathcal{A}=\{(j_{1},j_{2})|j_{1}+j_{2}\in\mathbb{Z}\}$ with $j_{1},j_{2}=0,1/2,1,\dots,k/2.$ Of these, $(k/2,k/2)$ is a transparent boson, which is sufficient to conclude that the decohered theory is a non-modular anyon theory.
@@
Thus, we can obtain a large family of imTOs by exposing the ground states of string-net models to local error channels, where the decohered code space generically retains logical information i.e., it is a decoherence-free subspace. The presence of non-trivial logical information (or a quantum memory) is encoded in the Wilson-loop algebra (equivalently, the $S$-matrix of $\mathcal{A}$). Since we have shown that the resulting imTOs can host transparent anyons, we obtain a partial classification of imTO in terms of braided fusion categories.
@@
V Higher-Form Symmetry and non-unitary exfoliation of Walker-Wang models
@@
Let us focus on the Toric code first for concreteness. The initial pure state TO density matrix trivially has a strong $\mathbb{Z}_{2}^{e}\times\mathbb{Z}_{2}^{m}$ 1-form symmetry. Working in the doubled space picture, we then see that the $\mathbb{Z}_{2}^{(0)}$ imTO resulting from $e$-decoherence still has a strong $\mathbb{Z}_{2}^{e}$ 1-form symmetry generated by Wilson loops associated to the deconfined, quantum excitation $e_{+}$, but only a weak $\mathbb{Z}_{2}^{m}$ 1-form symmetry, generated by $m_{+}m_{-}$. Indeed, in each of the examples we have studied, we see that decoherence has a non-trivial effect on the underlying strong 1-form symmetries of the parent TO. We can thus rephrase our results for generic Abelian TOs in the language of 1-form symmetry: when gauging out an anyon $a$ via decoherence, the resulting imTO density matrix retains strong 1-form symmetries for those symmetries generated by anyons which braid trivially with $a$, while the remaining 1-form symmetries are reduced to weak symmetries. In other words, only those 1-form symmetries which do not have a mixed anomaly with the 1-form symmetry generated by Wilson loops of $a$ remain as strong 1-form symmetries, while the remainder are reduced to weak symmetries. As noted previously, in the TSSC framework, the strong 1-form symmetries of the decohered state are manifest in Eq. (6), where the stabilizers may be viewed as closed Wilson loops for the deconfined anyons. This observation further reinforces the idea that the logical subsystem forms a decoherence free subspace under local noise that incoherently proliferates certain anyons, affecting only degrees of freedom in the gauge subsystem.
The 1-form symmetry structure of imTOs provides a useful language with which to characterize the stability of these states to local noise as well as their utility as memories  [^10] Here, we are specifically referring to stability under finite-depth local noise channels.. Regarding the former, the characterization of the decohered mixed-state in terms of strong 1-form symmetries immediately implies that these fixed point density matrices belong to an imTO phase, where each state in that phase shares the same 1-form symmetry structure and these states can be connected via strong 1-form symmetry preserving finite-depth LPQCs. Similar to gapped ground states of local Hamiltonians, we also expect stability up to some finite noise threshold under general local noise channels since LPQCs satisfy an area-law (i.e., satisfy a Lieb-Robinson bound) and cannot generate arbitrarily long-range correlations that destroy the state at infinitesimally small noise rates [65] (note however that a notion of a “gap” remains lacking in this context). Each strong 1-form symmetry, being equivalent to the presence of an anyon in the theory, moreover implies the existence of non-local operators commuting with the stabilizer group—the corresponding anyon Wilson loop along the non-contractible cycles of the torus.
@@
We may then employ the anomalies of the 1-form symmetries to characterize the structure of the code space. Specifically, if two 1-form symmetries have a mixed anomaly, they give rise to a pair of logical operators and hence a quantum memory, as in the $\mathbb{Z}_{4}^{(1)}$ TSSC. If a 1-form symmetry has no mixed anomalies but has a $\mathbb{Z}_{N}$ anomaly with $N>2$, its corresponding Wilson loops along the two cycles of the torus also yield logical operators and a quantum memory, as in the chiral semion TSSC. Finally, if a 1-form symmetry has no mixed anomalies and at most a $\mathbb{Z}_{2}$ anomaly (i.e. it is either a boson or fermion), its corresponding non-contractible Wilson loops only yield non-local stabilizers, thus yielding a classical memory. In passing we note that while our focus has been on the implications of strong symmetries, excitations charged under weak 1-form symmetries (hence these are confined), correspond to “fluxes” in the language of TSSCs [31, 36]. Higher form symmetries thus provide a convenient language with which to characterize the TSSC structure of imTOs.
@@
There is also a striking analogy between imTO and anomalous surface states of certain pure state 3D TOs, which suggests potential generalizations of our scheme to other intrinsically mixed-states. As we have emphasized throughout, imTO generally supports chiral and non-modular TO in a purely 2D system. In the context of local gapped Hamiltonians, such states naturally arise at the 2D surfaces of 3D topological orders, specifically those realized in the Walker-Wang (WW) models [38]. These are 3D exactly solvable lattice models which, given a potentially non-modular TO $\mathcal{C}$, realize $\mathcal{C}$ as its surface theory. In particular, if we consider a slab geometry with open boundary conditions in, say, the $z$ direction, one obtains $\mathcal{C}$ on the top surface and $\overline{\mathcal{C}}$ on the bottom surface. If $\mathcal{C}$ is modular, then the bulk has trivial topological order. Conversely, if $\mathcal{C}$ is non-modular, then the bulk is topologically ordered and supports both point-like and loop-like excitations, generated at the ends of Wilson lines and edges of Wilson surfaces, respectively, which braid non-trivially with each other. These loop-like excitations can be absorbed by the surfaces. Importantly, the transparent anyons in $\mathcal{C}$ also correspond to deconfined point-like excitations in the bulk, and so can freely move from the top $\mathcal{C}$ surface, into the bulk, and onto the bottom $\overline{\mathcal{C}}$ surface. Additionally, a “tube-like” Wilson surface stretching between a loop on the top surface and a loop on the bottom surface serves as a symmetry of the ground state, as the loop-like excitations are condensed on the surfaces.

### v3

@@ III.4 Three-Fermion from $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ Toric code
As a final remark, finite temperature mixed-states also provide simple instances of our general framework. For example, consider the $D=2,3,4$ Toric code at finite temperature [58, 12]. For $D=2$, any finite-T state corresponds to both the $e,m$ anyons being incoherently proliferated: the resulting state hosts no deconfined anyons and is hence trivial. Now, in $D=3$, the $e$ charges proliferate at any non-zero temperature, but below a critical $T_{c}$, the flux-loops of the 3D Toric code remain deconfined but are now transparent. This corresponds to a TSSC that does not encode any logical qubits in its logical subsystem but still has a non-trivial classical memory due to the transparent loops. Finally, for the 4D Toric code which has only loop-like excitations, there exist two critical temperatures: below the first, none of the excitations proliferate and the finite-T mixed-state is a TSSC that is equivalent to a topological stabilizer code (i.e., its gauge group is proportional to the stabilizer group). Hence, it represents finite-T topological order. Above the first, but below the second critical temperature, only one of the loop excitations proliferate and one obtains a TSSC with a classical memory. Above the second critical temperature, all anyons are condensed and the mixed-state is topologically trivial. Thus, prior results on topological order at finite temperature are straightforwardly incorporated into our general framework. We note that an infinite temperature state with quantum memory based on a subsystem code was previously proposed in Ref. [59], and in our framework, constitutes an imTO.
In general, the map from decohered density matrices to TSSCs conveniently allows one to use results regarding the latter to obtain a partial classification of the former. In particular, since Ref. [40] showed that any (non-modular) Abelian anyon theory can be realized by a TSSC, it immediately provides a partial classification of Abelian imTOs in terms of non-modular Abelian anyon theories.
@@ IV Decoherence as gauging out in general UMTCs
Thus, we can now define gauging out anyons in the same way as before, but in a more general context: starting with a UMTC $\mathcal{C}$ and a proper subset of anyons $\hat{\mathcal{A}}$ to be gauged out, the resulting anyon theory (the code space of the decohered theory) is given by those anyons in $\mathcal{C}$ which braid trivially with those in $\hat{\mathcal{A}}$. Moreover, if any anyons in $\hat{\mathcal{A}}$ are transparent in $\hat{\mathcal{A}}$, they remain transparent in $\mathcal{A}$, which will generically be a braided fusion category (without the modularity restriction). Formally, given a UMTC $\mathcal{C}$ and a proper subset of objects (anyons) $\hat{\mathcal{A}}$ (i.e., a full subcategory of $\mathcal{C}$), the anyon theory $\mathcal{A}$ that results upon gauging out $\hat{\mathcal{A}}$ is given by the centralizer $C_{C}(\hat{\mathcal{A}})$ of $\hat{\mathcal{A}}$ in $\mathcal{C}$:
$\mathcal{A}\equiv\{x\in\mathcal{C}|\mathcal{B}_{\theta}(x,y)=1\,\forall y\in\hat{\mathcal{A}}\},$ (17)
which is a braided fusion category (see Ref. [64]). One could in principle then generate another braided fusion category by gauging out anyons from $\mathcal{A}$ and generate a cascade of imTOs by iteratively gauging out anyons. As discussed in Section II, we define a density matrix with imTO as one where $\mathcal{A}$ is non-modular, even in the non-Abelian case.
We believe that this picture for obtaining braided fusion categories from parent UMTCs falls squarely within the general class of mixed topological quantum field theories (TQFTs) proposed by Zini and Wang in Ref. [65], but where the input to the parent 2+1D Turaev-Viro (TV) type TQFT is always modular. In our context, this restriction is physically motivated since we take as input the ground state of a local Hamiltonian (so the anyon theory is a UMTC) and then subject it to local noise. In fact, note that the doubled semion example we previously considered, in which the resulting mixed state supports the chiral semion TO, is presented as an example of a mixed TQFT in Ref. [65]. Specifically, in that case the input was the doubled semion UMTC and the output mixed TQFT was the chiral semion UMTC, where we can view the latter as the Reshetikhin-Turaev (RT) TQFT of the former. Ref. [65] described this as “tracing” out the anti-chiral degrees of freedom, which we believe corresponds to the process of gauging out presented above. This supports our claim that (unitary) non-modular braided fusion categories provide a partial classification of imTOs. More generally, if the parent theory is some doubled Chern-Simons topological quantum field theory (which admits a gapped boundary to vacuum), one might expect that local error channels will lead to the underlying chiral Chern-Simons TQFT in the decohered mixed-state—we hope to investigate this general procedure in the context of a continuum field theory description in a forthcoming work. For now, we consider some simple examples that go beyond Abelian anyon theories to show the generality of our framework.
@@ IV.1 $\mathbb{Z}_{2}^{(1)}$ TSSC from Chiral Ising UMTC
A physical Hamiltonian that supports a phase with chiral Ising TO is furnished by the Kitaev honeycomb model [50]. We consider gauging out the $\psi$ fermion: since $\sigma$ braids non-trivially with $\psi$, the resulting anyon theory describing the decohered state is simply given by $\{1,\psi\}$, which does not encode any quantum memory but still yields a classical memory and retains a well-defined fermionic excitation.
@@ IV.3 Non-modular imTOs from Doubled $SU(2)_{k}$ UMTC
As a final example, we can consider doubled SU(2)k string-net models, whose lattice models and short-string operators are given in Ref. [61]. Anyons in this theory are labeled by pairs $(j_{1},j_{2})$ where $j=0,1/2,1,\dots,k/2$. Let us now subject a ground state of this model to local errors that incoherently proliferate the anyon $(k/2,k/2)$ (which is a boson). In order to read off the resulting imTO in the decohered density matrix, after $(k/2,k/2)$ has been gauged out, we require the braiding relations of this theory. In particular, the braiding between an anyon $(j_{1},j_{2})$ and $(k/2,k/2)$ is given by $\mathcal{B}_{\theta}((j_{1},j_{2}),(k_{1},k_{2}))=(-1)^{2(j_{1}+j_{2})}$. Thus, the resulting imTO is characterized by the anyon theory $\mathcal{A}=\{(j_{1},j_{2})|j_{1}+j_{2}\in\mathbb{Z}\}$ with $j_{1},j_{2}=0,1/2,1,\dots,k/2.$ Of these, $(k/2,k/2)$ is a transparent boson, which is sufficient to conclude that the decohered theory is a non-modular anyon theory.
@@
Thus, we can obtain a large family of imTOs by exposing the ground states of string-net models to local error channels, where the decohered code space generically retains logical information i.e., it is a decoherence-free subspace. The presence of non-trivial logical information (or a quantum memory) is encoded in the Wilson-loop algebra (equivalently, the $S$-matrix of $\mathcal{A}$). Since we have shown that the resulting imTOs can host transparent anyons (corresponding to a non-modular anyon theory), we obtain a partial classification of imTO in terms of non-modular unitary braided fusion categories.
V Locally Detectable Anyons
Thus far, we have characterized the topological order exhibited by mixed states in terms of their anyon data $\mathcal{A}$ i.e., the set of anyons whose closed Wilson loops remain coherent after subjecting the original pure state to decoherence. In the Abelian case, these correspond to the stabilizer group and, as we will discuss below, in general these correspond to the set of strong 1-form symmetries respected by the mixed state. However, one could also characterize topological order in terms of the distinct anyon types which remain as locally detectable excitations outside the code space  [^10] We thank Michael Levin for a discussion prompting us to include this perspective.. Indeed, in pure state topological order, the detectable anyon types are generated by open Wilson lines, which are in one-to-one correspondence with the Wilson loops generating the ground state degeneracy on the torus. In contrast, while in the present context there is no Hamiltonian and hence no notion of an excitation gap (see however Ref. [67]), one may still identify states (outside the code space) with local errors as “excited states” of the mixed state. As we now show, if the TO characterizing the mixed state $\mathcal{A}$ is modular, then $\mathcal{A}$ also describes the set of locally detectable anyon types. However, if $\mathcal{A}$ is non-modular, then $\mathcal{A}$ is no longer in one-to-one correspondence with the set of locally detectable anyons. The physical meaning of this will be made clearer in the following section, when we identify imTO as corresponding to the surface of a Walker-Wang model.
Let us consider a mixed state $\rho$ obtained via decoherence of an anyon $a$ in a parent TO $\mathcal{C}$. Then, from the preceding discussion, $\rho$ is characterized by the (potentially non-modular) TO $\mathcal{A}$ obtained by gauging out $\hat{\mathcal{A}}=\{a^{m}\}$ in $\mathcal{C}$. Now, let $c$ be an anyon in the parent TO $\mathcal{C}$ and $W_{x_{1},x_{2}}^{c}$ the Wilson line operator creating $c$ and its conjugate $\overline{c}$ at well-separated points $x_{1}$ and $x_{2}$. Likewise, let $W_{\Gamma}^{d}$ be the operator creating a Wilson loop of $d\in\mathcal{C}$ along the contractible cycle $\Gamma$. Then we will call
$\rho_{c}=W_{x_{1},x_{2}}^{c}\rho(W_{x_{1},x_{2}}^{c})^{\dagger}$ (18)
an “excited state” relative to $\rho$ if we can detect the presence of $c$ via braiding with some anyon $d$. That is, we wish to compute
$\mathrm{Tr}[W_{\Gamma}^{d}\rho_{c}]=\mathrm{Tr}[W_{\Gamma}^{d}W_{x_{1},x_{2}}^{c}\rho(W_{x_{1},x_{2}}^{c})^{\dagger}]=\mathcal{B}_{\theta}(d,c)\mathrm{Tr}[W_{\Gamma}^{d}\rho]\,,$ (19)
where $\Gamma$ encloses only, say, the point $x_{1}$ and we have used the braiding between anyons $c$ and $d$. In order for this expression to be non-zero, we require $W_{\Gamma}^{d}$ to be a stabilizer of $\rho$ and hence $d$ must be in $\mathcal{A}$. Note that there is no such restriction on $c$. However, $c$ is an excitation if and only if $\mathcal{B}_{\theta}(d,c)\neq 1$. Note that this means that transparent anyons in $\mathcal{A}$ are not locally detectable anyon types while all of the opaque anyons in $\mathcal{A}$ represent genuine quantum excitations. In particular, any two anyons in the parent theory $\mathcal{C}$ which remain detectable after decoherence and differ by fusion with a transparent anyon become indistinguishable as excitations. Thus, given an imTO $\mathcal{A}$ that results from incoherently proliferating anyons in a UMTC $\mathcal{C}$, the set of locally detectable anyons is given by
=$\mathcal{L}=\{c\in\mathcal{C}|\mathcal{B}_{\theta}(c,d)\neq 1\text{ for any }d\in\mathcal{A}\}\,,$ (20)
with the identification that $r\sim r\otimes t$, where $r\in\mathcal{L}$ and $t\in\mathcal{T}$, the set of transparent anyons. The mathematical structure underlying $\mathcal{L}$ remains mysterious; note for instance that it need not even be closed under fusion!
In the doubled space picture, we thus see that the locally detectable anyon types in the original Hilbert space correspond to anyons of the form $(c_{+},c_{-})$, while the observables (namely the Wilson loops which detect other anyons via braiding and hence the stabilizers) are in correspondence with anyons of the form $(d_{+},1)$.
Let us understand the structure of the set of locally detectable anyons by way of a few examples. First consider a parent TO that can be written as a product of two modular TOs, $\mathcal{C}=\mathcal{A}\boxtimes\hat{\mathcal{A}}$, and subject it to decoherence that gauges out $\hat{\mathcal{A}}$ – the DS theory discussed above is one such example. Since, by definition, every $d\in\mathcal{A}$ braids trivially with every $c\in\hat{\mathcal{A}}$, the only locally detectable and hence genuine quantum excitations, are labelled by the anyons in $\mathcal{A}$. Thus, for a modular theory, the set of locally detectable anyons types is in one-to-one correspondence with the braided fusion category $\mathcal{A}$ characterizing the mixed state TO.
This correspondence does not hold for non-modular imTO. Indeed, let us consider the simple example of the Toric code subjected to $e$ deocherence, yielding the $\mathbb{Z}_{2}^{(0)}$ mixed state discussed above. While the $\mathbb{Z}_{2}^{(0)}$ is characterized by the anyon content $\{1,e\}$, the $e$ anyon does not label a genuine quantum excitation, as it is a transparent anyon. This is trivially seen from the fact that $Z_{i}\rho Z_{i}=\rho$. In contrast, the original $m$ and $f$ anyons of the parent Toric code do exist as genuine quantum excitations, as they may be detected via expectation values of $W^{e}_{\Gamma}$ Wilson loops  [^11] Note that while the $e$ anyons do not exist as genuine quantum excitations, we can nevertheless use braiding of $e$ to detect other anyons.. Moreover, the $m$ and $f$ anyons are identified as excitations in the decohered theory, as there are no stabilizers that can distinguish their spin.
While for a generic imTO, the anyon theory (equivalently, the braided fusion category) $\mathcal{A}$ describing the mixed state $\rho$ does not label the set of locally detectable anyon excitations, the latter is still fully determined by the former. As such, the set $\mathcal{L}$ provides a finer characterization of the imTO, which should be a feature of all states within the same imTO phase and not simply fixed point states. We elaborate on this in the following section in the language of anomalies between strong and/or weak symmetries. This framework also clarifies how the set of locally detectable anyon types can be determined simply by knowing the symmetries of $\rho$ and their associated anomalies.
@@ V Locally Detectable Anyons
VI Higher-Form Symmetry and non-unitary exfoliation of Walker-Wang models
@@
Let us focus on the Toric code first for concreteness. The initial pure state TO density matrix trivially has a strong $\mathbb{Z}_{2}^{e}\times\mathbb{Z}_{2}^{m}$ 1-form symmetry. Working in the doubled space picture, we then see that the $\mathbb{Z}_{2}^{(0)}$ imTO resulting from $e$-decoherence still has a strong $\mathbb{Z}_{2}^{e}$ 1-form symmetry generated by Wilson loops associated to the anyon $e_{+}$, but only a weak $\mathbb{Z}_{2}^{m}$ 1-form symmetry, generated by $m_{+}m_{-}$. Indeed, in each of the examples we have studied, we see that decoherence has a non-trivial effect on the underlying strong 1-form symmetries of the parent TO. We can thus rephrase our results for generic Abelian TOs in the language of 1-form symmetry: when gauging out an anyon $a$ via decoherence, the resulting imTO density matrix retains strong 1-form symmetries for those symmetries generated by anyons which braid trivially with $a$, while the remaining 1-form symmetries are reduced to weak symmetries. In other words, only those 1-form symmetries which do not have a mixed anomaly with the 1-form symmetry generated by Wilson loops of $a$ remain as strong 1-form symmetries, while the remainder are reduced to weak symmetries. As noted previously, in the TSSC framework, the strong 1-form symmetries of the decohered state are manifest in Eq. (6), where the stabilizers may be viewed as closed Wilson loops for the deconfined anyons. This observation further reinforces the idea that the logical subsystem forms a decoherence free subspace under local noise that incoherently proliferates certain anyons, namely those affecting only degrees of freedom in the gauge subsystem.
@@
In this language, we can hence rephrase our characterization of imTO as follows: a density matrix $\rho$ exhibits imTO if its set of strong 1-form symmetries form a non-modular anyon theory – that is, its set of strong 1-form symmetries cannot be consistently realized in the ground state of a gapped, local 2D Hamiltonian. The 1-form symmetry structure of imTOs provides a useful language for characterizing the utility of these states as quantum memories. Note that each strong 1-form symmetry implies the existence of non-local operators commuting with the stabilizer group—the corresponding anyon Wilson loop along the non-contractible cycles of the torus. We may then employ the anomalies of the 1-form symmetries to characterize the structure of the code space. Specifically, if two strong 1-form symmetries have a mixed anomaly, they give rise to a pair of logical operators and hence a quantum memory, as in the $\mathbb{Z}_{4}^{(1)}$ TSSC. If a strong 1-form symmetry has no mixed anomalies but has a $\mathbb{Z}_{N}$ anomaly with $N>2$, its corresponding Wilson loops along the two cycles of the torus also yield logical operators and a quantum memory, as in the chiral semion TSSC. Finally, if a strong 1-form symmetry has no mixed anomalies and at most a $\mathbb{Z}_{2}$ anomaly (i.e. it is either a boson or fermion), its corresponding non-contractible Wilson loops only yield non-local stabilizers, thus yielding a classical memory. Higher form symmetries thus provide a convenient language with which to characterize the TSSC structure of Abelian imTOs.
Indeed, the mixed anomalies between strong and weak 1-form symmetries play a central role in determining the set of locally detectable anyons $\mathcal{L}$ in the imTO. In particular, any strong 1-form symmetry of a state $\rho$ that has a $\mathbb{Z_{N}}$ self-anomaly (with $N>2$) is associated with a locally detectable anyon. If any strong 1-form symmetries have a mixed anomaly, they too correspond to anyons in the imTO. One can think of such anomalous strong 1-form symmetries as corresponding to $\mathcal{A}\backslash\mathcal{T}$ i.e., the anyons in the (generically non-modular) braided fusion category $\mathcal{A}$ minus the set of transparent anyons $\mathcal{T}$ in $\mathcal{A}$. Finally, weak 1-form symmetries which have a mixed anomaly with a strong 1-form symmetry also correspond to locally detectable anyons (see Sec. V); while the transparent anyons do not have any self anomalies (with $\mathbb{Z}_{N>2}$) or mixed anomalies with other strong symmetries, they can have mixed anomalies with weak 1-form symmetries and therefore influence the structure of $\mathcal{L}$. Hence, the anomaly structure of the strong and weak symmetries of a short-range correlated 2D mixed state $\rho$ provides a detailed characterization of the corresponding imTO.
@@
There is also a striking analogy between imTO and anomalous surface states of certain pure state 3D TOs, which suggests potential generalizations of our scheme to other intrinsically mixed-states. This also allows us to generalize the discussion from the preceding paragraph to the non-Abelian case. As we have emphasized throughout, imTO generally supports chiral and non-modular TO in a purely 2D system. In the context of local gapped Hamiltonians, such states naturally arise at the 2D surfaces of 3D topological orders, specifically those realized in the Walker-Wang (WW) models [42]. These are 3D exactly solvable lattice models which, given a potentially non-modular TO $\mathcal{A}$, realize $\mathcal{A}$ as its surface theory. In particular, if we consider a slab geometry with open boundary conditions in, say, the $z$ direction, one obtains $\mathcal{A}$ on the top surface and $\bar{\mathcal{A}}$ on the bottom surface. If $\mathcal{A}$ is modular, then the bulk has trivial topological order. Conversely, if $\mathcal{A}$ is non-modular, then the bulk is topologically ordered and supports both point-like and loop-like excitations, generated at the ends of Wilson lines and edges of Wilson surfaces, respectively, which braid non-trivially with each other. These loop-like excitations can be absorbed by the surfaces. Importantly, the transparent anyons in $\mathcal{A}$ also correspond to deconfined point-like excitations in the bulk, and so can freely move from the top $\mathcal{A}$ surface, into the bulk, and onto the bottom $\bar{\mathcal{A}}$ surface. Additionally, a “tube-like” Wilson surface stretching between a loop on the top surface and a loop on the bottom surface serves as a symmetry of the ground state, as the loop-like excitations are condensed on the surfaces.